In a triangle, one angle is 6 degrees more than the second angle. The third angle is 18 degrees more than the second angle. Find the measure of each angle.
Answer by checkley77(12844)

F=FIRST ANGLE.
S=SECOND ANGLE.
T=THIRD ANGLE.
F=S+6
T=S+18
F+S+T=180
(S+6)+S+(S+18)=180
3S+24=180
3S=180-24
3X=156
S=156/3
S=52 DEGREES.
F=52+6=58 DEGREES.
T=52+18=70 DEGREES.
PROOF:
52+58+70=180
180=180